### v2

# Matrix-Driven Instant Review: Confident Detection and Reconstruction of LLM Plagiarism on PC

Abstract
In recent years, concerns about intellectual property (IP) in large language models (LLMs) have grown significantly. Plagiarizing other LLMs (through direct weight copying, upcycling, pruning, or continual pretraining) and claiming authorship without properly attributing to the original license, is a serious misconduct that can lead to significant financial and reputational harm to the original developers. However, existing methods for detecting LLM plagiarism fall short in key areas. They fail to accurately reconstruct weight correspondences, lack the ability to compute statistical significance measures such as $p$-values, and may mistakenly flag models trained on similar data as being related. To address these limitations, we propose Matrix-Driven Instant Review (MDIR), a novel method that leverages matrix analysis and Large Deviation Theory. MDIR achieves accurate reconstruction of weight relationships, provides rigorous $p$-value estimation, and focuses exclusively on weight similarity without requiring full model inference. Experimental results demonstrate that MDIR reliably detects plagiarism even after extensive transformations, such as random permutations and continual pretraining with trillions of tokens. Moreover, all detections can be performed on a single PC within an hour, making MDIR both efficient and accessible.
1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has resulted in a proliferation of models trained on vast amounts of data, achieving impressive performance across various tasks. However, this progress has also raised concerns about the potential misuse of model weights, such as direct copying, upcycling [52, 19], pruning [25, 1], and continual pretraining, followed by claims of authorship without proper attribution to the original developers. These practices, commonly referred to as "plagiarizing" other LLMs, severely undermine the integrity of the research community and erode public trust. The complexity and size of LLMs make it particularly challenging to detect such plagiarism in a timely and efficient manner.
Existing methods for detecting model similarity can be broadly classified into two main categories: retrieval-based methods and representation-based methods.
Retrieval-based Methods. Retrieval-based methods [49] rely on vendors designing specific key-value pairs ${(k_{i},v_{i})}_{i=1,\cdots,N}$ for language model watermarking. These $(k_{i},v_{i})$ pairs are typically generated randomly (e.g., (k=’396f4b37...’, v=’586a99’)) to ensure they rarely appear in conventional conversational scenarios. During pretraining, vendors inject these key-value pairs into the training data to maximize the conditional probability $p_{\theta}(v_{i}|k_{i})$. If a downstream model $\theta^{\prime}$ exhibits significantly higher $p_{\theta^{\prime}}(v_{i}|k_{i})$ than a random guessing baseline (or lower perplexity; e.g., perplexity =16=16 for random hexadecimal digits), vendors may confidently claim the origin of the downstream model. This method has proven effective in identifying model plagiarism, such as the case of Llama3-V plagiarizing MiniCPM-o v2.6 [33], where rare oracle bone inscriptions were used as keys and their corresponding modern Chinese characters as values. However, this approach requires access to vendor-specific keys or prompts, which introduces significant limitations. In practice, it is often impractical for users to obtain or know the internal training data of vendors.
Representation-based Methods. Representation-based methods, such as REEF (Representation Encoding Fingerprint) [55], HuRef (Human-Readable Fingerprint) [53], and Intrinsic Fingerprint [19], analyze similarities between internal representations of language models. Similarities can be examined both in terms of weights and activation distributions under the same input. An ideal similarity measure should remain robust to simple attacks, such as dimension permutation, scaling, or orthogonal transformations. Moreover, this measure should remain stable throughout training, even after continual pretraining with trillions of tokens.
To the best of our knowledge, these methods primarily focus on proving similarity through "fingerprints" but lack the ability to identify, trace, or reconstruct the specific process of model plagiarism.
Our proposed method overcomes these limitations by operating at the matrix level, directly computing similarity without requiring access to vendor-specific data. Furthermore, we provide deeper insights by identifying exact relationships, such as permutations and transformations, during model copying. We also introduce statistical significance via $p$-values, a critical component for making robust statistical inferences that is absent in prior research. Our advancements address a key gap in current methodologies, offering both efficiency and interpretability in detecting model similarity.
Our method is grounded in a strong mathematical foundation, leveraging Singular Value Decomposition (SVD) and polar decomposition to analyze model matrices. We further utilize Large Deviation Theory (LDT) and random matrix theory to estimate $p$-values. These mathematical tools provide a rigorous framework for detecting similarities at a fundamental level, making it exceedingly difficult for cheaters to bypass our detection. Cheaters attempting to copy models would face significant challenges in obfuscating this watermark.
Our method also democratizes the model verification process. Without requiring access to vendor-specific prompts or specialized hardware, anyone with a standard PC can participate in the verification process.
2 Preliminaries
Our work primarily builds upon the foundation of matrix analysis. To avoid excessive technicality or verbosity, we only introduce the key tools and properties used in this paper, potentially omitting or abbreviating proofs. Interested readers are encouraged to refer to [21] for a comprehensive overview of matrix analysis.
2.1 Matrix Analysis: Singular Value and Polar Decompositions
Singular Value Decomposition. The singular value decomposition (SVD) of a matrix $A\in\mathbb{R}^{m\times n}$ is given by
$A=USV^{T},$
where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices, and $S\in\mathbb{R}^{m\times n}$ is a diagonal matrix with non-negative singular values on the diagonal:
$\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}>0\quad(r=\mathrm{rank}(A)).$
We denote $\sigma_{i}(A)$ as the $i$-th singular value of $A$.
Polar Decomposition. The polar decomposition of $A$ has two common but distinct forms: the left decomposition $A=PW$ and the right decomposition $A=WQ$, where $P=(AA^{T})^{1/2}$ and $Q=(A^{T}A)^{1/2}$ are symmetric positive semidefinite matrices, and $W\in\mathbb{R}^{m\times n}$ (with orthonormal columns) is shared between both decompositions. We define the orthogonal part of $A$ as $\mathrm{Ortho}(A):=W$. When $A$ is full-rank, the orthogonal part of $A$ is unique.
Connection to SVD. The orthogonal part $W$ in the polar decomposition can be obtained from the SVD:
$W=UV^{T},$
where $U$ and $V$ are derived from the SVD of $A$. The symmetric factors satisfy $P=USU^{T}$ and $Q=VSV^{T}$. However, when $A$ is not invertible (with many singular values either exactly zero or close to zero), the computation of $W=UV^{T}$ becomes ill-conditioned.
SVD and Spectral Calculus. Assume $A\in\mathbb{R}^{m\times n}$ has SVD $A=USV^{T}$. Then:
$AA^{T}=U(SS^{T})U^{T},\quad A^{T}A=V(S^{T}S)V^{T}.$
For any polynomial function $f\in\mathbb{R}[x]$ and $g(x)=xf(x^{2})$, we have:
$f(AA^{T})A$
$=Uf(SS^{T})U^{T}USV^{T}=Uf(SS^{T})SV^{T}$
$=Ug(S)V^{T}$
$=USf(S^{T}S)V^{T}=USV^{T}Vf(S^{T}S)V^{T}$
$=Af(A^{T}A).$
In fact, $g(S)$ can represent any odd polynomial. For any odd function $G$, we may find a polynomial $f$ such that $f(x)$ agrees with $G(x)/x$ on all nonzero diagonal entries of $SS^{T}$. In particular, setting $G(x)=\mathrm{sign}(x)$, this spectral calculus yields the most “sensible” orthogonal part of $A$. We adopt the methods in [3] for a fast and relatively accurate implementation.
Trace Maximization Property. The orthogonal part $W$ solves the optimization problem:
$\max_{\{X|X^{T}X=I\}}\left(\operatorname{Tr}(AX^{T})\right).$
The maximum value is the sum of all the singular values of $A$, i.e., $\operatorname{Tr}(P)=\sum_{i=1}^{r}\sigma_{i}$.
Orthogonal Invariance of Singular Values. The singular values $\{\sigma_{i}\}$ of $A$ are invariant under both left and right orthogonal transformations:
$\sigma_{i}(UAV)=\sigma_{i}(A),\quad\forall U,V\text{ orthogonal and }i=1,\cdots,\min(m,n).$
Thus, any function depending on these singular values remains invariant under orthogonal transformations. Examples include:
  • Frobenius Norm: $\lVert A\rVert_{F}=\sqrt{\sum_{i}\sigma_{i}^{2}(A)}$;
  • Spectral Norm: $\lVert A\rVert_{S}=\sigma_{1}(A)$;
  • Ky Fan $k$-Norm: $\lVert A\rVert_{KF}=\sum_{i=1}^{k}\sigma_{i}(A)$;
  • Schatten $p$-Norm [37]: $\lVert A\rVert_{p}=\sqrt[p]{\sum_{i}\sigma_{i}^{p}(A)}$.
Any combination of these functions also remains invariant under orthogonal transformations. These functions may serve as preliminary indicators for detecting model similarity.
Orthogonal Matrices and RMSNorms Commute. For any orthogonal matrix $U$, we have:
$\mathrm{RMSNorm}(x)U=\mathrm{RMSNorm}(xU),$
for any nonzero vector $x\in\mathbb{R}^{n}$ (row vector). Moreover, all transformations satisfying this property $(\mathrm{RMSNorm}(\cdot))U=\mathrm{RMSNorm}((\cdot)U)$ are orthogonal transformations.
To prove this, note that $\mathrm{RMSNorm}(y)$ is always a constant multiple ($\sqrt{n}$) of a unit vector when $y\neq 0$. Thus, $U$ maps all unit vectors to unit vectors. By linearity, this implies $\lVert xU\rVert=\lVert x\rVert$ and $xUU^{T}x^{T}=xx^{T}$. Taking $x$ over all eigenspaces of $UU^{T}$, we see that $1$ is the only eigenvalue of $UU^{T}$. Hence, $UU^{T}=\bm{1}_{n}$, and $U$ is orthogonal.
2.2 Large Deviation Theory
Our research extensively involves random orthogonal matrices, particularly focusing on traces of such matrices. To obtain statistically meaningful $p$-values, traditional statistical $p$-tests become useless for our case, due to the interdependence of the elements within an orthogonal matrix. Instead, we rely on large deviation theory for deriving the $p$-value. For a comprehensive exposition of this theory, please refer to [11] and [4].
3 Methods
3.1 Problem Statement
Model plagiarism refers to the act of reusing part or all of another model’s weights without proper acknowledgment as required by the original model’s usage protocol. In this paper, we focus solely on weight reuse and do not address similarities arising from training data selection.
Given two large language models $A$ and $B$, with their parameters denoted as $\theta_{A}$ and $\theta_{B}$, we aim to determine whether $A$ and $B$ exhibit a relationship in their weights, based solely on the statistical properties of $\theta_{A}$ and $\theta_{B}$. The weight relationships we consider include, but are not limited to, the following cases:
  • Fine-tuning: Various kinds of SFT and RL included;
  • Continual Pretraining: Training the model with more data in the general domain, sometimes as much as trillions of tokens;
  • Upcycling [52, 19]: Continual pretraining with a larger model with weights initialized from a smaller base model;
  • Pruning [25, 1]: Removing certain channels or neurons of a base model to obtain a smaller model;
  • Transformation: Should include permutations and even general orthogonal/unitary transformations;
  • Comprehensive Obfuscation: A combination of all these cases above.
This task can be formulated as a binary classification problem $\Psi$, where the inputs are $(\theta_{A},\theta_{B})$ and the output is $\Psi(\theta_{A},\theta_{B})\in\{0,1\}$. Here, $1$ indicates that the two models are related, while 0 indicates they are unrelated.
Due to the vast number of parameters (on the order of billions) and the scarcity of samples, it is impractical to construct a machine learning framework that reliably determines whether two models share a weight relationship. Any learning-based approach to this task would likely suffer from extreme overfitting without producing trustworthy signals. To address this challenge, we leverage tools from statistics and random matrix theory.
Our algorithm focuses exclusively on detecting relationships between the weights $\theta_{A}$ and $\theta_{B}$. Consequently, all arguments presented below remain valid if $\theta_{A}$ and $\theta_{B}$ are swapped. For convenience, we may assume that model $A$ is the victim (reference model) and model $B$ is the suspect. However, it is important to note that our algorithm does not distinguish between the suspect and the victim. The identification of the suspect, victim, and plagiarism should be determined using additional evidence, particularly temporal and societal clues.
3.2 Architectural Assumptions
For simplicity, we assume that both model $A$ and model $B$ adopt a decoder-only Transformer architecture [47, 35], with word embeddings and unembeddings, Grouped Query Attention (GQA) [2], MLP layers with up and down projections [41], and RMSNorm layers [54].
Most other architectures can be analyzed within our framework, as GQA represents the most prevalent form of attention mechanism in modern Transformers. Both Multi-Head Attention (with an expansion rate of $1$) [47] and Multi-Query Attention (with one key-value head per layer) [40] can be treated as special cases of GQA.
While our method may also apply to other architectures, such as linear RNN variants, a specialized analysis tailored to each architecture is required for accurate results.
We consider the GQA architecture depicted in Figure 1, which is illustrated using Penrose tensor notation.
Figure 1: A Penrose notation of Grouped Query Attention (top) and an illustration of an equivalent model under certain transformations. Bias terms are omitted, and RMSNorms are not explicitly shown.
3.3 Equivalent Transformations
The following analysis has been partially derived in prior literature [53], but with significant gaps. For instance, previous work often omits critical components such as Rotary Positional Embeddings (RoPE), attention heads, RMSNorms, and other architectural details.
Assume that Model $A$ admits the weights depicted in the upper subfigure of Figure 1, and Model $B$ is equivalent to $A$ under certain transformations:
$\theta_{A}$ $=\{Q,K,V,O\}$
$\theta_{B}$ $=\{Q^{\prime},K^{\prime},V^{\prime},O^{\prime}\}.$
The linear weight transformation from $\theta_{A}$ to $\theta_{B}$ may take the following form:
$Q^{\prime}$ $=U_{Q}QW_{Q},\quad K^{\prime}=U_{K}KW_{K},\quad V^{\prime}=U_{V}VW_{V},\quad O^{\prime}=W_{O}^{-1}OU_{O}^{-1},$
where $U_{Q},U_{K},U_{V},U_{O}$ and $W_{Q},W_{K},W_{V},W_{O}$ are transformation matrices applied to the original weights. These matrices represent modifications introduced during plagiarism or model adaptation.
We refer to $U_{Q},U_{K},U_{V},U_{O}$ as outer transformations, which typically correspond to operations such as rotations, permutations, or scaled orthogonal transformations. Since both $U_{Q},U_{K},U_{V}$ and $Q,K,V$ operate on normalized vectors:
$\mathrm{RMSNorm}(x^{\prime})U_{\square}=\mathrm{RMSNorm}(x),\quad\square\in\{Q,K,V\},$
this implies that $U_{Q}=U_{K}=U_{V}$, and they are all orthogonal matrices. In the context of Lie groups, we denote $U_{Q}=U_{K}=U_{V}\in\mathrm{O}(\mathrm{EmbDim})$. In contrast, $U_{O}$ can be less restrictive and may include a scaling factor: $U_{O}=\lambda U_{Q}$ ($\lambda\neq 0$). In Lie group notation, $U_{O}\in\mathrm{O}(\mathrm{EmbDim})\times\mathbb{R}$.
For the inner transformations $W_{Q},W_{K},W_{V},W_{O}$, the situation becomes more complex due to the presence of attention heads and nonlinear transformations (e.g., Softmax) across channels. Not all orthogonal transformations are permissible for inner transformations.
From the Penrose diagram in Figure 1, we propose a possible set of transformations for the inner matrices $W_{Q},W_{K},W_{V},W_{O}$. While we cannot prove that this set encompasses all possible transformations, it provides a plausible characterization:
$W_{Q}$ $=\mu\cdot P_{1}\otimes P_{2}\otimes S,$
$W_{K}$ $=\mu^{-1}\cdot P_{1}\otimes S,$
$V^{\prime}$ $=P_{1}\otimes H_{v},$
$W_{O}$ $=P_{1}\otimes P_{2}\otimes H_{v},$
where $\otimes$ denotes the tensor product of matrices, $\mu\neq 0$ is a scalar, $P_{1}\in\mathrm{Perm}(\mathrm{NumHeads})$ is a permutation matrix over $\mathrm{NumHeads}$ channels, $P_{2}\in\mathrm{Perm}(\mathrm{QueriesPerHead})$ is a permutation matrix over $\mathrm{QueriesPerHead}$ queries, $S\in\mathrm{diag}(\pm 1,\cdots,\pm 1)\in\mathrm{M}_{\mathrm{HeadDim}}(\mathbb{R})$, and $H_{v}\in\mathrm{GL}(\mathrm{HeadDim},\mathbb{R})$ is an arbitrary invertible matrix.
This set does not account for all possible configurations. For example, $S$ could be replaced by an arbitrary orthogonal matrix. However, such substitutions might obfuscate the implementation of Rotary Positional Embeddings (RoPE) [43], potentially flagging the model for plagiarism. To ensure compatibility with both QK-norm [20] and RoPE, we restrict $S$ to diagonal matrices with entries $\pm 1$:
$S\in\mathrm{diag}(\pm 1,\cdots,\pm 1).$
In practice, these transformation matrices may all reduce to identity matrices, particularly when the victim and suspect models are just identical.
3.3.1 Transformations with Training and Noise Injection
After applying all equivalent transformations, Model $B$ still inherits its policy entirely from Model $A$, i.e., $\pi(B)=\pi(A)$. As a result, $B$ remains easily detectable via retrieval-based watermarks, KL divergence, or simple output comparisons.
To obscure these patterns, the model stealer might train the model on a large number of tokens. In such cases, our formulas can be adjusted as follows:
$Q^{\prime}$ $=U_{Q}QW_{Q}+N_{Q},$
$K^{\prime}$ $=U_{K}KW_{K}+N_{K},$
$V^{\prime}$ $=U_{V}VW_{V}+N_{V},$
$O^{\prime}$ $=W_{O}^{-1}OU_{O}^{-1}+N_{O},$
where $N_{Q},N_{K},N_{V},N_{O}$ are noise matrices. We assume that the noise has a very small root-mean-square (RMS) magnitude compared to the original matrices $Q,K,V,O$.
3.3.2 Pruning and Upcycling
Without loss of generality, we may assume that Model $B$ is a pruned version of Model $A$. This formulation is analogous to Model $A$ being upcycled from Model $B$.
For simplicity, we provide a possible relationship:
$Q^{\prime}$ $=U_{Q}QW_{Q}+N_{Q},$
$K^{\prime}$ $=U_{K}KW_{K}+N_{K},$
$V^{\prime}$ $=U_{V}VW_{V}+N_{V},$
$O^{\prime}$ $=W_{O}OU_{O}^{T}+N_{O},$
where $U_{Q}=U_{K}=U_{V}=U_{O}=U$ is the same partial isometry (i.e., $UU^{T}=\bm{1}$); $W_{Q}$ and $W_{K}$ are also partial isometries satisfying $W_{Q}^{T}W_{Q}=\bm{1}$ and $W_{K}^{T}W_{K}=\bm{1}$. The possible forms of $W_{V}$ and $W_{O}$ involving pruning can be highly complex. We leave their exact characterization for future work.
3.4 Preliminary Examinations
To effectively identify model plagiarism, we must provide sufficient evidence and even reconstruct the entire fabrication process. To achieve this goal, we must systematically solve all permutations and transformations involved.
Previous research [53] correctly identified the following covariant relationships:
$Q^{\prime}K^{\prime T}$ $\approx UQK^{T}U^{T},$
$V^{\prime}O^{\prime}$ $\approx UVOU^{T},$
where $U$ is an orthogonal transformation matrix. Since there is only one matrix $U$ to be solved, it is feasible to estimate $U$ by minimizing the error:
$\min_{U\in\mathrm{O}(\mathrm{EmbDim})}\left(\lVert Q^{\prime}K^{\prime T}-UQK^{T}U^{T}\rVert_{F}^{2}+\lVert V^{\prime}O^{\prime}-UVOU^{T}\rVert_{F}^{2}\right).$
Because $U$ represents the same transformation across all layers, we can extend this optimization to all layers:
$\min_{U\in\mathrm{O}(\mathrm{EmbDim})}\sum_{l=1}^{L}\left(\lVert Q^{\prime}_{l}K^{\prime T}_{l}-UQ_{l}K^{T}_{l}U^{T}\rVert_{F}^{2}+\lVert V^{\prime}_{l}O^{\prime}_{l}-UV_{l}O_{l}U^{T}\rVert_{F}^{2}\right),$
or equivalently,
$\max_{U\in\mathrm{O}(\mathrm{EmbDim})}\sum_{l=1}^{L}\left(2\mathrm{Tr}(UQ_{l}K^{T}_{l}U^{T}K^{\prime}_{l}Q^{\prime T}_{l})+2\mathrm{Tr}(UV_{l}O_{l}U^{T}O^{\prime T}_{l}V^{\prime T}_{l})\right).$
Despite systematic approaches for maximizing a single term of $\mathrm{Tr}(UAU^{T}B)$ [8], this optimization problem involving multiple such terms is challenging and may not admit a unique solution. For example: If $U$ is a solution, then U-U is also a solution. In extreme cases, if all $Q_{l}K^{T}_{l}$, $Q^{\prime}_{l}K^{\prime T}_{l}$, $V^{\prime}_{l}O^{\prime}_{l}$, and $V_{l}O_{l}$ are identity matrices, then any $U\in\mathrm{O}(\mathrm{EmbDim})$ satisfies the equations.
3.5 Breaking the Shell: Vocabulary Embeddings
Let $E\in\mathbb{R}^{\mathrm{VocabSize}\times\mathrm{EmbDim}}$ denote the vocabulary embedding matrix of model $A$, and $E^{\prime}\in\mathbb{R}^{\mathrm{VocabSize}\times\mathrm{EmbDim}}$ for model $B$. Following the conventions established in previous discussions, we have:
$E^{\prime}=EU^{T}+N_{E},$
where $U\in\mathbb{R}^{\mathrm{EmbDim}\times\mathrm{EmbDim}}$ is an orthogonal matrix, and $N_{E}$ represents the additional perturbation introduced by training or noise injection.
To minimize the difference between $E^{\prime}$ and $EX^{T}$, we solve the following optimization problem:
$\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\lVert E^{\prime}-EX^{T}\rVert_{F}^{2}=\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\left\langle E^{\prime}-EX^{T},E^{\prime}-EX^{T}\right\rangle_{F}.$
Expanding the Frobenius norm yields:
$\arg\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\lVert E^{\prime}-EX^{T}\rVert_{F}^{2}$ $=\arg\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\left(\lVert E^{\prime}\rVert_{F}^{2}+\lVert E\rVert_{F}^{2}-2\left\langle EX^{T},E^{\prime}\right\rangle_{F}\right)$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\left\langle EX^{T},E^{\prime}\right\rangle_{F}$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left(EX^{T}E^{\prime T}\right)$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left((E^{\prime T}E)X^{T}\right).$
We denote $\tilde{U}$ as the solution to this optimization problem, which corresponds to the orthogonal part of $(E^{\prime T}E)$. Note that $\tilde{U}$ is not the ground truth of $U$, but rather a close approximation. If $\tilde{U}$ is sufficiently close to a permutation matrix, we can identify a permutation matrix $P\in\mathrm{Perm}(\mathrm{EmbDim})$ via:
$P=\arg\max\operatorname{Tr}(P\tilde{U}^{T}).$
This problem is equivalent to solving maximum bipartite matching or the linear sum assignment problem [9], which can be computed in $O(n^{3})$ time using the Hungarian algorithm. For large matrices (e.g., $18432\times 18432$), solving this exactly may take approximately one hour on a CPU. To reduce computational cost, we propose a faster heuristic: compute the positions of the maximal element in each row and verify whether they form a valid permutation.
A unique determination of $\tilde{U}$ requires $(E^{\prime T}E)$ to be non-degenerate and full-rank, which necessitates the vocabulary size to satisfy $\mathrm{VocabSize}\geq\mathrm{EmbDim}$. This condition is easily satisfied, as most recent tokenizers have $\mathrm{VocabSize}\geq 3\times 10^{4}$.
3.5.1 When the Tokenizer Is Changed
If the model stealer attempts to obfuscate the embeddings by using a different tokenizer, the situation becomes more complex. However, we expect a significant overlap in tokens between model $A$’s and model $B$’s tokenizer, including:
  • Single-byte ASCII characters (the smallest unit in byte-pair encoding [38]);
  • Common English words such as am, is and take;
  • Accented letters like é.
These tokens collectively constitute a large subset of the vocabulary, and we anticipate the size of their intersection to exceed $1\times 10^{4}$.
Drawing from Saussure’s theory of linguistics [36], tokens are arbitrary, but their contextual usage systematically reveals their meaning. Based on this assumption, even after reinitialization, the embeddings of shared tokens, trained on billions of words, should gradually align.
Let $\mathcal{C}$ denote the set of all common tokens. We estimate $\tilde{U}$ as:
$\tilde{U}=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left(\left(E^{\prime}[\mathcal{C},:]^{T}E[\mathcal{C},:]\right)X^{T}\right).$
Thus, $\tilde{U}$ corresponds to the orthogonal part in the polar decomposition of $(E^{\prime}[\mathcal{C},:]^{T}E[\mathcal{C},:])$.
3.6 Estimating the $p$-Value
We aim to identify a permutation matrix $P$ from $\tilde{U}$ as $P=\arg\max\operatorname{Tr}(P\tilde{U}^{T})$. However, we need a statistical criterion to determine whether our identification is successful. While it may be reasonable to visually inspect plots of $P$ and $\tilde{U}$, the value of $\operatorname{Tr}(P\tilde{U}^{T})$ itself often provides strong evidence. In some experiments, this value can reach as high as $0.2\cdot\mathrm{EmbDim}$ to $0.99\cdot\mathrm{EmbDim}$.
Null Hypothesis. Our null hypothesis assumes that models $A$ and $B$ are not homologous, and there is no apparent similarity between their weights. Under this assumption, the probability measure $\mathrm{d}\mathbb{P}$ should be uniform across all admissible transformations. For instance, for the vocabulary embedding matrix $E$ of model $A$, its density satisfies:
$E\mathrm{d}\mathbb{P}=(EG)\mathrm{d}\mathbb{P},\quad\forall G\in\mathrm{O}(n),$
where $G$ is an orthogonal matrix. Consequently, the distribution of $\tilde{U}$, as the orthogonal part of $(E^{\prime T}E)$, satisfies:
$\tilde{U}\mathrm{d}\mathbb{P}=(\tilde{U}G)\mathrm{d}\mathbb{P}.$
This implies that $\tilde{U}$ is uniformly distributed over $\mathrm{O}(n)$ according to the Haar measure.
Now, fix $P_{0}$ as an arbitrary permutation matrix. We estimate the probability of $\operatorname{Tr}(P_{0}\tilde{U}^{T})\geq c$, denoted as:
$f(c):=\mathbb{P}\left[\operatorname{Tr}(P_{0}\tilde{U}^{T})\geq c\right].$
Since $\tilde{U}^{T}$ is uniformly distributed under the Haar measure, we have:
$(P_{0}\tilde{U}^{T})\mathrm{d}\mathbb{P}=(P_{0}P_{0}^{T}\tilde{U}^{T})\mathrm{d}\mathbb{P}=\tilde{U}^{T}\mathrm{d}\mathbb{P}.$
Thus:
$f(c)=\mathbb{P}\left[\operatorname{Tr}(\tilde{U})\geq c\right].$
However, our permutation $P$ maximizes $\operatorname{Tr}(P\tilde{U}^{T})$. With $n!$ possible permutations, only the one maximizing $\operatorname{Tr}(P\tilde{U}^{T})$ is chosen. This implies that the $p$-value satisfies:
$p$ $=\max_{P\in\mathrm{Perm}(n)}\mathbb{P}\left[\operatorname{Tr}(P\tilde{U})\geq c\right]$
$=\mathbb{P}\left[\exists P\in\mathrm{Perm}(n),\ \operatorname{Tr}(P\tilde{U})\geq c\right]$
$\leq\sum_{P_{0}\in\mathrm{Perm}(n)}\mathbb{P}\left[\operatorname{Tr}(P_{0}\tilde{U})\geq c\right]$
$=n!\cdot f(c).$
This result indicates that we can estimate the $p$-value based on the evaluation of $f(c)$.
[…]
$\mathrm{TV}(f-(1-\Phi))\leq\exp(-\alpha n),\quad\text{for some }\alpha>0,$
where $\Phi(x)$ is the cumulative distribution function of the standard normal distribution.
It is known that $(1-\Phi)(x)$ has the following asymptotic behavior for large $x$:
$(1-\Phi)(x)\approx\frac{1}{\sqrt{2\pi}x}\exp\left(-\frac{x^{2}}{2}\right),\quad(x\gg 1).$
However, when both $n$ and $x$ are large, $\exp(-\alpha n)$ remains significantly larger than $\frac{1}{\sqrt{2\pi}x}\exp\left(-\frac{x^{2}}{2}\right)$. Thus, it is inappropriate to use $(1-\Phi)$ as the asymptotics of $f(c)$.
To study the tail behavior of $f(c)$, we leverage tools from Large Deviation Theory. Since $n$ is typically large and the embedding dimension $n$ is usually an even number in most models, we assume $n=2m$ without loss of generality. Additionally, we assume $\det(\tilde{U})=1$, or equivalently $\tilde{U}\in\mathrm{SO}(2m)$, since this introduces only an infinitesimal difference in the thermodynamic limit ($n\to\infty$; see Appendix A). These assumptions simplify our analysis while preserving accuracy.
Theorem 1.
Let $A\in\mathrm{SO}(2m)$ be uniformly distributed according to the normalized Haar probability measure, and let $0<r\leq 1/2$. The probability
$P(r,m)=\mathbb{P}\left[\frac{1}{2m}\operatorname{Tr}(A)\geq r\right]$
satisfies the following large deviation principle:
$\lim_{m\to\infty}\frac{-\log P(r,m)}{2m^{2}r^{2}}=1,$
or equivalently,
$P(r,m)\asymp\exp(-m^{2}I(r)),$
where
$I(r)=2r^{2}$
is the good rate function.
For $r>1/2$, a faster decay rate can be achieved: $I(r)>2r^{2}$.
For the proof, please refer to Appendix A.
Based on this result, our $p$-value can now be estimated as follows:
$p$ $\leq n!\cdot f(c)$
$\approx n!\cdot P(c/n,m)$
$\asymp n!\cdot\exp(-2m^{2}c^{2}/n^{2})$
$=n!\cdot\exp\left(-\frac{c^{2}}{2}\right).$
However, this estimation remains coarse and imprecise. To provide definitive statistical evidence of model similarity, we recommend adopting a threshold of at least ten standard deviations ($10\sigma$), corresponding to a $p$-value of less than $2\times 10^{-23}$. Achieving this criterion is feasible in the context of billion-parameter-scale models, as even tiniest similarities will yield extremely significant results.
Refining the $p$-value estimation further would require advanced mathematical tools and techniques, which are beyond the scope of this research. This topic may warrant a separate investigation in future work.
3.7 Solving the Rest
If the significance of the $p$-value has already been determined from the previous calculations, this step is entirely optional. However, if the $p$-value is not significant at this stage, we cannot yet rule out the possibility of plagiarism. This may occur when the outer transformation is a general orthogonal matrix. Additionally, interested readers may wish to determine the exact relationship between two models, potentially with different architectures.
Since $P=\arg\max\operatorname{Tr}(P\tilde{U}^{T})$, we proceed as follows: If $P$ is reliably identified as a permutation matrix ($p<2\times 10^{-23}$), set $U:=P$; otherwise, set $U:=\tilde{U}$.
Solving the Transformations in the Attention Module. We now solve for the inner transformations $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$ based on the heuristic $U$.
Our objectives are:
$\min_{W_{Q}\text{ orthogonal}}\lVert UQW_{Q}-Q^{\prime}\rVert_{F}^{2},\quad\min_{W_{K}\text{ orthogonal}}\lVert UKW_{K}-K^{\prime}\rVert_{F}^{2},$
for $Q$ and $K$. By the trace maximization property, the solutions for $W_{Q}$ and $W_{K}$ are given by:
$W_{Q}=\lambda\mathrm{Ortho}(Q^{T}U^{T}Q^{\prime}),\quad W_{K}=\mu\mathrm{Ortho}(K^{T}U^{T}K^{\prime}),$
where the scaling coefficients $\lambda$ and $\mu$ are computed as:
$\lambda=\frac{\lVert Q^{\prime}\rVert_{F}}{\lVert Q\rVert_{F}},\quad\mu=\frac{\lVert K^{\prime}\rVert_{F}}{\lVert K\rVert_{F}}.$
For $W_{V}$ and $W_{O}$, if the model stealer uses an orthogonal transformation, we apply the same method to compute $\mathrm{Ortho}(V^{T}U^{T}V^{\prime})$ and $\mathrm{Ortho}(OU^{T}O^{\prime T})$. However, since $W_{V}$ and $W_{O}$ may involve general invertible transformations, solving the general case without additional assumptions is challenging. We leave this problem for future work.
Solving the MLP. At this point, there is only one matrix left to solve: the permutation of intermediate neurons, denoted by $P$. The solution is given as follows:
$U_{X}$ $=\mathrm{Ortho}(X^{T}U^{T}X^{\prime}),\quad X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\};$
$P$ $=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(P\left(U_{\mathrm{Gate}}+U_{\mathrm{Up}}+U_{\mathrm{Down}}\right)^{T}\right).$
Typically, we expect the three solutions $\arg\max\operatorname{Tr}\left(PU_{X}^{T}\right)$, for $X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\}$, to yield the same permutation. However, computing the orthonormal part for intermediate matrices (which often have 14,000–20,000 rows) is computationally expensive. Adding the three terms together would triple the computation cost. If the noise level is tolerable, we may simply select one of them:
$P=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(PU_{\mathrm{Up}}^{T}\right).$
Similarly, the $p$-value for permuted MLPs can be estimated as:
$x$ $=\operatorname{Tr}\left(PU_{\mathrm{Up}}^{T}\right),$
$p_{\mathrm{MLP}}$ $=\mathrm{IntermediateDim}!\cdot\exp(-x^{2}/2).$
However, this is not an accurate estimation for the MLP, as similar features may gradually emerge during training, and the correspondence between neurons may also arise independently.
3.8 Overall Algorithm
We summarize our algorithm as follows:
Algorithm 1 Computing the $p$-value: $\mathrm{PValue}(P,U,d)$
Input: Permutation matrix $P$, orthogonal matrix $U$, dimension $d$
Compute $p:=d!\cdot\exp(-(\operatorname{Tr}(PU^{T}))^{2}/2)$
Output: $p$
Algorithm 2 MDIR (Matrix-Driven Instant Review)
Input: Model $A$, Model $B$
Initialize: Threshold $p_{0}=2\times 10^{-23}$, relation flag $r=\mathsf{False}$
$L:=$ Number of layers in the model
$E:=$ Embedding matrix of AA
$E^{\prime}:=$ Embedding matrix of $B$
$\mathcal{C}:=$ Set of common vocabulary tokens between $A$ and $B$
Compute $\tilde{U}:=\mathrm{Ortho}(E[\mathcal{C},:]^{T}E^{\prime}[\mathcal{C},:])$ via polar decomposition
$P:=\arg\max_{P\in\mathrm{Perm}(\mathrm{EmbDim})}\operatorname{Tr}(P\tilde{U}^{T})$
Compute $p:=\mathrm{PValue}(P,\tilde{U},\mathrm{EmbDim})$
if $p<p_{0}$ then
  Set $r:=\mathsf{True}$
  Set $U:=P$
else
  Set $U:=\tilde{U}$
end if
Yield: Transformation matrix $U$, $p$-value $p$
for Layer $i\in[L]$ do
  Extract attention weights $Q,K,V,O$ from layer $i$ of AA
  Extract attention weights $Q^{\prime},K^{\prime},V^{\prime},O^{\prime}$ from layer $i$ of $B$
  for $X\in\{Q,K,V\}$ do
   Compute $W_{X}:=\mathrm{Ortho}(X^{T}U^{T}X^{\prime})\cdot\frac{\lVert X^{\prime}\rVert_{F}}{\lVert X\rVert_{F}}$
  end for
  Extract MLP weights $\mathrm{Gate},\mathrm{Up},\mathrm{Down}$ from layer $i$ of AA
  Extract MLP weights $\mathrm{Gate}^{\prime},\mathrm{Up}^{\prime},\mathrm{Down}^{\prime}$ from layer $i$ of $B$
  Compute $U_{X}:=\mathrm{Ortho}(X^{T}U^{T}X^{\prime})$ for $X\in\{\mathrm{Up}\}$ (or $X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\}$)
  $P:=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}(PU_{\mathrm{Up}}^{T})$
  Compute $p:=L\cdot\mathrm{PValue}(P,U_{\mathrm{Up}},\mathrm{IntermediateDim})$
  if $p<p_{0}$ then
   Yield: Transformation matrix $U_{\mathrm{Up}}$, $p$-value $p$
   Set $r:=\mathsf{True}$
  end if
end for
Output: Relation flag $r$
4 Experiments and Case Studies
We conducted a series of experiments to evaluate the effectiveness of our MDIR method in identifying similarities between models. Considering file size constraints, for each matrix, we plotted the first $512\times 512$ submatrix. The complete plots of the entire matrices can be produced by the accompanying code.
4.1 Level 1: Identifying Instruction-tuned Versions and Avoiding False Positives for Unrelated Models
This section examines official instruction-tuned versions of models in comparison to their base counterparts:
  • Qwen2.5-0.5B and Qwen2.5-0.5B-Instruct [34]: Figure LABEL:fig:qwen25_0b5
  • Meta-Llama-3.1-8B and Meta-Llama-3.1-8B-Instruct [28]: Figure LABEL:fig:llama31_8b
These pairs of models are known to be unrelated, even through they share similar architectures:
  • Meta-Llama-3.1-8B and Qwen3-8B-Base [51]: Figure LABEL:fig:llama3_qwen3
  • DeepSeek-V3-Base [10] and Kimi-K2-Instruct [30]: Figure LABEL:fig:ds_kimi
4.2 Level 2: Identifying Upcycling Models
These models are known as upcycled from their predecessors, according to their technical reports:
  • Qwen1.5-1.8B [44] and Qwen1.5-MoE-A2.7B [45]: Figure LABEL:fig:qwen15_18
  • Qwen2-7B and Qwen2-57B-A14B [50]: Figure LABEL:fig:qwen2_57
4.3 Level 3: Identifying Continual Pretraining of Pruned and Distilled Models
These models are pruned versions of their predecessors and have subsequently undergone continual pretraining:
  • Meta-Llama-3.1-8B and Meta-Llama-3.2-1B [1]: Figure LABEL:fig:llama31_llama32_1b
  • Meta-Llama-3.1-8B and Meta-Llama-3.2-3B: Figure LABEL:fig:llama31_llama32_3b
We observe that the large values, shown by red pixels, form irregular sloped lines in the relation matrices. Each row (from the smaller model) corresponds to a single large value associated with one channel from the larger model. This pattern indicates that the smaller model is a pruned version of the larger model.
Our computed $p$-value is heavily overestimated, and the actual $p$-value should be orders of magnitude smaller. Since the channels of the smaller model are arranged in an orderly manner, there are only $\mathbf{C}(n,m)$ admissible configurations for such a pattern, not $n!$. Unfortunately, we have not derived the correct large deviation principle for non-square semi-orthogonal matrices like in this example. Therefore, we decide to retain this over-estimated $p$-value, and leave the refinement of the analysis to future work.
4.4 Level 4: Identifying Continual Pretraining of Non-Transformer Models with Architectural Modifications
In addition to Transformer-based models, our work also detects relationships between models of non-Transformer architectures:
  • RWKV-7-World-0.4B [32] and RWKV-5-Eagle-0.4B [31]: Figure LABEL:fig:rwkv57
  • RWKV-7-World-2.9B and RWKV-6-Finch-3B: Figure LABEL:fig:rwkv67
The RWKV series (RWKV-5, 6 and 7) uses a head size of $64\times 64$, which differs from a traditional $128\times 128$ head size commonly used in most modern Transformers. This $64\times 64$ block pattern is clearly visible from Figure LABEL:fig:rwkv57(b). However, with significant architectural changes, the diagonal patterns become less distinct and more challenging to identify.
4.5 Level 5: Identifying Plagiarism in Continual Pretraining of Upscaled Models with Comprehensive Obfuscation
A recent debate has emerged regarding whether the Pangu-Pro-MoE model [6, 5] is derived from Qwen2.5-14B. Our analysis provides strong evidence supporting this claim with a high level of confidence.
  • Pangu-Pro-MoE [6] and Qwen2.5-14B: Figure LABEL:fig:pangu_qwen
To solve the relationship between the embedding channels, we identified 101,486 tokens in common between the tokenizer of Pangu-Pro-MoE (which contains 153,376 tokens) and that of Qwen2.5-14B (with 151,665 tokens). The initial step in solving the relation matrix from these vocabulary embeddings of these overlapping tokens yielded a trace of 3862, corresponding to a $p$-value less than $10^{-3,000,000}$. Subsequent analyses consistently produced similarly low $p$-values, indicating a significant similarity in the embeddings, attention modules, and shared expert/MLP modules between the two models. As illustrated in Figure LABEL:fig:pangu_qwen, the red dots in the nearly white background, though barely visible, reveal that these matrices are nearly permutation matrices. Taking into account the temporal and contextual factors, such as the fact that Pangu-Pro-MoE was released approximately nine months after Qwen2.5-14B, we conclude that Pangu-Pro-MoE reuses weights from Qwen2.5-14B. This finding raises concerns about the validity and reliability of the technical report of Pangu-Pro-MoE [5].
4.6 Ablation: MDIR Only Detects Weight Relevance, Not Training Data
We demonstrate that our MDIR method exclusively detects relevance in weights, rather than training data. To validate this claim, we conducted an ablation experiment by initializing two models with different random seeds and training each on two distinct datasets, resulting in a total of four models.
The datasets used were:
  • DCLM subsample [24] (the first 100 files of shard 0 [29], containing 12.05 billion tokens), and
  • OpenWebMath-ProX [56] (4.61 billion tokens [15], abbreviated as OWM hereafter).
Both datasets were tokenized using the GPT-NeoX tokenizer [7].
[…]
Figure LABEL:fig:seed reveals clear diagonal patterns for models initialized with the same seed, indicating strong weight similarity due to shared initialization. In contrast, Figure LABEL:fig:samedata shows varying brightness blocks for solved attention relation matrices but no significant outliers, suggesting that models trained on the same dataset may develop similar attention features without substantial weight correlation.
4.7 Comparative Analysis
REEF Method May Produce False Positives. We compare our method against REEF (Representation Encoding Fingerprint) [55] using the following models:
  • Meta-Llama-3.1-8B vs. Meta-Llama-3.2-3B: True positive (Figure LABEL:fig:llama31_llama32_3b),
  • Meta-Llama-3.1-8B vs. Qwen2.5-7B: False positive (Figure LABEL:fig:reef).
The results indicate a similarity score exceeding $0.8$ for the first few layers (e.g., layer 2 of Meta-Llama-3.1-8B vs. layer 2 of Qwen2.5-7B), despite the fact that the Llama 3.1 and Qwen2.5 series are entirely unrelated.
This phenomenon may arise because both models, after training on tens of trillions of tokens, gradually develop similar representations, which can be detected by CKA (Central Kernel Alignment). In contrast, our MDIR method is specifically designed to detect weight similarity, not representation similarity, and safely returns negative results in such cases (Figure LABEL:fig:comparison).
5 Conclusion
We introduced MDIR, a novel method for LLM weight similarity and plagiarism detection, leveraging robust mathematical tools and providing statistically meaningful results. Our method does not require full model inference and can be executed on a single PC or even without a GPU, democratizing the verification process. Moreover, our method detects only weight similarity but not feature similarity, making it more specific and targeted than previous techniques. This reduces the likelihood of false positives and enhances reliability.
5.1 Limitations
Numerical Precision. All of our discussions above assume infinite precision; however, practically, our method relies on matrix polar decomposition, which is an ill-posed problem when the matrix is low-rank. Moreover, since most neural networks are trained using low-precision formats (fp16, bf16, or even fp8), this challenges our assumption that transformation matrices are uniformly distributed across the Haar measure. In practice, we also observed approximately a 1% difference when computing the polar part under fp32 and fp64.
Extreme $p$-values. Our method produces $p$-values on the order of $10^{-3,000,000}$, far exceeding $p$-values observable in traditional statistical settings. This is reasonable given the scale of modern models, which often contain billions of parameters, compared to the hundreds or thousands of samples typically analyzed in conventional statistics. However, this may raise concerns when combined with numerical precision issues. For example, assuming a typical error of 2%, and if unfortunately $25\sigma$ of error deviates from the infinite-precision ground truth, it could literally halve the trace and diminish the $p$-value significance obtained from the trace. Therefore, the actual $p$-value, though still highly significant statistically, is much larger than that reported by our method.
5.2 Future Work
Potential Ways of Evading Detection. Although not observed in our experiments, there might be ways to evade our detection. We hypothesize that MDIR could be evaded through additional training with larger learning rates, especially for vocabulary embeddings. We leave this for future work.
More Accurate Estimation of $p$-values. We addressed the $p$-value for the orthogonal case (square matrices) using large deviation theory. However, accurate estimation of $p$-values for the semi-orthogonal case (non-square matrices, also known as partial isometries and Stiefel manifolds) remains an open challenge. This is more common because the embedding and intermediate dimensions often differ between two arbitrary models. We also leave this important refinement for future work.
6 Ethics Statement
In developing our MDIR framework, we deeply respect intellectual property rights and adhere to ethical AI practices. We strongly discourage any form of plagiarism or unauthorized fabrication of large language models. Such actions not only undermine innovation but also erode trust within the AI community.
Our work aims to promote the democratization, transparency, and responsibility of AI technologies. Therefore, we firmly oppose the misuse of our framework to conceal plagiarism in LLMs.
Appendix A Background and Proof of Theorem 1
Before we delve into the proof, it is worth discussing the challenges we encountered during the process.
We are dealing with the Circular Real Ensemble (CRE, named after [39]), which has a formulation distinct from traditional circular ensembles, such as Circular Unitary Ensembles (CUE) or Circular Orthogonal Ensembles (COE) [13]. These ensembles play important roles in both random matrix theory and condensed matter physics.
One of the problems most similar to ours is the Gross-Witten ensemble [18, 16], which concerns the large deviations from the typical value of $\mathrm{Re}\operatorname{Tr}(U)$ as $N\to\infty$ (where $U\in\mathrm{U}(N)$ is a random unitary matrix).
Unitary matrices possess many elegant properties. One of them is the Cayley transformation, defined as $\phi(z):=\mathrm{i}(1+z)/(1-z)$, which maps the punctured unit circle $\mathbb{T}\setminus\{1\}$ to the real line $\mathbb{R}$. When the Cayley transformation is applied to a unitary matrix, it sends unit eigenvalues to the real line, thereby producing a Hermitian matrix (i.e., $\phi(U)$ is Hermitian for $U\in\mathrm{U}(N)$). However, under the Cayley transformation, an orthogonal matrix is transformed into a purely imaginary, anti-symmetric matrix, which is of little interest. In the field of deep learning, complex values rarely appear, and we shall focus on the Circular Real Ensemble in our proof.
The circular ensemble is closely related to thermodynamics and large deviation theory. We refer to Mehta’s book Random Matrices [27], and a comprehensive introduction can be found in Chapter 12 of it.
A.1 Proof of Theorem 1
Proof.
If $A$ is uniformly distributed according to the Haar measure in $\mathrm{SO}(2m)$ (a manifold of real dimension m(2m1)m(2m-1)), all eigenvalues of $A$ lie on the unit circle, and complex eigenvalues form paired conjugates, possibly accompanied by several +1+1 and 1-1.
When $\det A=1$, the product of complex eigenvalues yields +1+1, and there are almost surely no 1-1 or +1+1 eigenvalues.
Denote by $\{e^{i\theta_{k}},e^{-i\theta_{k}}:k=1,2,\cdots,m\}$ the eigenvalues of $A$. It is a classical result (see [48, 17]) that the phases $(\theta_{k})_{k}$ obey the distribution characterized by the following probability density:
$p(\theta_{1},\cdots,\theta_{m})\mathrm{d}\theta_{1}\cdots\mathrm{d}\theta_{m}=C\prod_{1\leq k<j\leq m}(\cos\theta_{k}-\cos\theta_{j})^{2}\mathrm{d}\theta_{1}\cdots\mathrm{d}\theta_{m},$
and the trace $\operatorname{Tr}(A)$ is the sum of all eigenvalues:
$\operatorname{Tr}(A)=2\sum_{i=1}^{m}\cos(\theta_{i}).$
By substitution of variables, let $t_{i}=\cos(\theta_{i})\in[-1,1]$, and $\mathrm{d}t_{i}/\sqrt{1-t_{i}^{2}}=\mathrm{d}\theta_{i}$. We study the substituted distribution:
$p(t_{1},\cdots,t_{m})\mathrm{d}t_{1}\cdots\mathrm{d}t_{m}=C^{\prime}\prod_{1\leq k<j\leq m}(t_{k}-t_{j})^{2}\cdot\prod_{1\leq i\leq m}(1-t_{i}^{2})^{-1/2}\mathrm{d}t_{1}\cdots\mathrm{d}t_{m}.$
Taking the logarithm, we have
$\log p(t_{1},\cdots,t_{m})=\sum_{1\leq k<j\leq m}\left(-2\log|t_{k}-t_{j}|\right)+\sum_{1\leq i\leq m}\left(\frac{\log(1-t_{i}^{2})}{2}\right)+C_{0}.$
This has a clear thermodynamical interpretation: Consider $m$ interacting particles located on the interval [1,1][-1,1], with coordinates $t_{1},\cdots,t_{m}$. The energy $E(t_{1},\cdots,t_{m})$ is the sum of the following two kinds of potentials:
  1. 1.
    Repelling force: for particles $k$ and $j$, their interaction potential is $2\log|t_{k}-t_{j}|$, meaning that they repel each other according to the 2-dimensional Coulomb law;
  2. 2.
    External field: particles are attracted to the boundary points 1-1 and +1+1, with the potential $\sum_{1\leq i\leq m}\left(\frac{\log(1-t_{i}^{2})}{2}\right)$.
Inspired by Chapter 12 of [27] and also [46], we define the Canonical Ensemble of these $m$ particles as follows, which admits a partition function:
$Z(t_{1},\cdots,t_{m})$ $=\int_{-1}^{1}\cdots\int_{-1}^{1}\exp({-\beta E(t_{1},\cdots,t_{m})})\mathrm{d}t_{1}\cdots\mathrm{d}t_{m},$
where
$E(t_{1},\cdots,t_{m})$ $=\sum_{1\leq k<j\leq m}\left(-2\log|t_{k}-t_{j}|\right)+\sum_{1\leq i\leq m}\left(\frac{\log(1-t_{i}^{2})}{2}\right)$
$=-\log p(t_{1},\cdots,t_{m})-C_{0}.$
We set the thermodynamic beta to be $\beta=1/(k_{B}T)=1$ because we do not study temperature changes.
We are interested in the probability:
$P(r,m)=\mathbb{P}\left[\frac{1}{2m}\operatorname{Tr}(A)\geq r\right].$
Using the representation $\operatorname{Tr}(A)=2\sum_{k=1}^{m}\cos(\theta_{k})=2\sum_{k=1}^{m}t_{k}$, this becomes:
$P(r,m)=\mathbb{P}\left[\frac{1}{m}\sum_{k=1}^{m}t_{k}\geq r\right].$
Let
$\mu_{m}=\frac{1}{m}\sum_{k=1}^{m}\delta_{t_{k}}$
be the empirical measure of the eigenvalue distribution. In the thermodynamic limit $m\to\infty$, $\mu_{m}$ should converge weakly to an equilibrium measure $\mu$ that minimizes the free energy functional $F(\mu)$.
We now solve the exact form of $F(\mu)$. We refer to Theorem 2.1 of the paper [14], where we are dealing with the case $\theta=1$, $\kappa=1$, and $w_{m}(x)=(1-x^{2})^{1/(2m)}$, where $w_{m}(x)\to 1$ in the thermodynamic limit $m\to\infty$. This suggests that the interaction term dominates, and the external field term is negligible when $m$ is large.
The form of $F(\mu)$ is therefore given by
$F(\mu)=\iint_{[-1,1]^{2}}\left(\log\frac{1}{|x-y|}\right)\mathrm{d}\mu(x)\mathrm{d}\mu(y),$
with the rate function
$\tilde{I}(\mu)=\iint_{[-1,1]^{2}}\left(\log\frac{1}{|x-y|}\right)\mathrm{d}\mu(x)\mathrm{d}\mu(y)-c,$
where $c=\inf_{\mu}F(\mu)$. Note that the rate function $\tilde{I}(\mu)$ is defined for the probability measure $\mu$, not yet ready for our $I(r)$.
From Corollary 2.2 of the paper [14], it suffices to solve the following two variational problems:
$\inf_{\mu}F(\mu);\quad\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}F(\mu).$
To solve these two problems, we parametrize $\mu$ using Chebyshev polynomials, with the additional assumption that $\mu$ is absolutely continuous with respect to the Lebesgue measure $\mathrm{d}x$ and normalized Chebyshev measure $\mathrm{d}x/(\pi\sqrt{1-x^{2}})$.
We suppose that $\mu$ is parametrized by the following series:
$\mu(x)=\frac{1}{\pi\sqrt{1-x^{2}}}\sum_{i=0}^{\infty}a_{i}T_{i}(x),$
where $T_{i}$ is the $i$-th Chebyshev polynomial of the first kind, and $F(\mu)$ now becomes a quadratic form of these coefficients $\{a_{i}\}$. An additional constraint must not be overlooked: $\mu$ is a probability measure, and
$\int_{-1}^{1}\mathrm{d}\mu=\frac{T_{0}(x)}{\pi\sqrt{1-x^{2}}}\sum_{i=0}^{\infty}a_{i}T_{i}(x)\mathrm{d}x=a_{0}=1.$
It is known that $\log|x-y|$ has the Chebyshev expansion [42, 26]:
$\log|x-y|=-\log 2-\sum_{n=1}^{\infty}\frac{2}{n}T_{n}(x)T_{n}(y),$
so
$F(\mu)$ $=\iint_{[-1,1]^{2}}\left(\log 2+\sum_{n=1}^{\infty}\frac{2}{n}T_{n}(x)T_{n}(y)\right)\frac{1}{\pi^{2}\sqrt{1-x^{2}}\sqrt{1-y^{2}}}$
$\quad\quad\quad\left(\sum_{i=0}^{\infty}a_{i}T_{i}(x)\right)\left(\sum_{i=0}^{\infty}a_{i}T_{i}(y)\right)\mathrm{d}x\mathrm{d}y$
$=\log 2\cdot\iint_{[-1,1]^{2}}\frac{1}{\pi^{2}\sqrt{1-x^{2}}\sqrt{1-y^{2}}}\mathrm{d}x\mathrm{d}y$
$\quad+\iint_{[-1,1]^{2}}\left(\sum_{n=1}^{\infty}\frac{2}{n}T_{n}(x)T_{n}(y)\right)\frac{1}{\pi^{2}\sqrt{1-x^{2}}\sqrt{1-y^{2}}}$
$\quad\quad\quad\quad\left(\sum_{i=1}^{\infty}a_{i}T_{i}(x)\right)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(y)\right)\mathrm{d}x\mathrm{d}y.$
We recall the orthogonal relation of Chebyshev polynomials,
$\int_{-1}^{1}\frac{T_{a}(x)T_{b}(x)}{\pi\sqrt{1-x^{2}}}\mathrm{d}x=\left\{\begin{aligned} &1,\quad&a=b=0;\\ &\frac{1}{2},\quad&a=b\neq 0;\\ &0,\quad&a\neq b.\end{aligned}\right$.
Using the orthogonal relation, we further simplify $F(\mu)$ as
$F(\mu)=\log 2+\sum_{i=1}^{\infty}\frac{a_{i}^{2}}{2i}.$
Therefore, the problem $\inf_{\mu}F(\mu)$ admits a simple solution: $a_{i}=0$ for every $i\geq 1$, and
$\mathrm{d}\mu_{0}$ $=\frac{\mathrm{d}x}{\pi\sqrt{1-x^{2}}};$
$F(\mu_{0})$ $=\log 2.$
Now we solve the equilibrium measure under one additional constraint $\int x\mathrm{d}\mu\geq r$:
$\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}F(\mu).$
Since
$\int_{-1}^{1}\mathrm{d}\mu=\frac{T_{1}(x)}{\pi\sqrt{1-x^{2}}}\sum_{i=0}^{\infty}a_{i}T_{i}(x)\mathrm{d}x=\frac{a_{1}}{2}=r,$
this suggests that the equilibrium measure $\mu_{r}$ is
$\mu_{r}(x)=\frac{1+2rx}{\pi\sqrt{1-x^{2}}}\quad(r\leq 1/2),$
and
$\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}F(\mu)=F(\mu_{r})=\log 2+2r^{2}.$
Now we are close to completion: our rate function $\tilde{I}(\mu)$ is computed as
$\tilde{I}(\mu)$ $=F(\mu)-c=F(\mu)-F(\mu_{0})=F(\mu)-\log 2.$
We conclude that the rate is
$\exp\left(-m^{2}\cdot\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}\tilde{I}(\mu)\right)$
$=\exp\left(-m^{2}\cdot(\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}\tilde{F}(\mu)-F(\mu_{0}))\right)$
$=\exp\left(-m^{2}\cdot 2r^{2}\right),$
which shows a good rate function for $r$: $I(r)=2r^{2}$.
The function $\mu_{r}$ induces a measure if and only if $r\leq 1/2$. When $r>1/2$, the function $\mu_{r}$ no longer induces a measure, as it violates positivity near x=1x=-1. This suggests that our theoretical bound, $\log 2+(2r)^{2}$, cannot be attained here (To satisfy the positivity of $\mu$, the coefficients $a_{2},a_{3},\cdots$ cannot all be zero simultaneously).
Overall, we have
$\inf_{\mu:\ \int x\mathrm{d}\mu\geq r}\tilde{I}(\mu)\geq(2r)^{2}$
and
[…]
A.2 Additional Notes on the $\mathrm{O}(2m)\setminus\mathrm{SO}(2m)$ Case
When $A\in\mathrm{O}(2m)\setminus\mathrm{SO}(2m)$, $\det(A)=-1$. There are two eigenvalues fixed at +1+1 and 1-1, respectively.
Our formula becomes:
$p(t_{1},\cdots,t_{m-1})\mathrm{d}t_{1}\cdots\mathrm{d}t_{m-1}=C^{\prime}\prod_{1\leq k<j\leq m-1}(t_{k}-t_{j})^{2}\cdot\prod_{1\leq i\leq m-1}(1-t_{i}^{2})^{+1/2}\mathrm{d}t_{1}\cdots\mathrm{d}t_{m-1}.$
Taking the logarithm, we have
$\log p(t_{1},\cdots,t_{m-1})=\sum_{1\leq k<j\leq m-1}\left(-2\log|t_{k}-t_{j}|\right)-\sum_{1\leq i\leq m-1}\left(\frac{\log(1-t_{i}^{2})}{2}\right)+C_{0}.$
The only differences from above are that $m$ is replaced by m1m-1 and the external field term is negated. Both changes are negligible in the thermodynamic limit $m\to\infty$. Thus, we obtain the same conclusion for the rate function.
A.3 The Exact Form of $I(r)$
We conjecture that the exact form of $I(r)$ is:
$\left\{\begin{aligned} I(r)&=2r^{2},\quad&\text{for }r\in\left(0,\frac{1}{2}\right];\\ I(r)&=\frac{1}{2}-\log 2-\log(1-r),\quad&\text{for }r\in\left[\frac{1}{2},1\right).\end{aligned}\right$.
Like the Gross-Witten Ensemble [18], there is a third-order phase transition near $r=1/2$.